OpenStudy (anonymous):

simplify (3√x^7) ÷ (√ x^3)

OpenStudy (anonymous):

You will need the identity \(\frac{\sqrt{a}}{\sqrt{b}} = \sqrt{\frac{a}{b}}\). \[ \frac{3\sqrt{x^7}}{\sqrt{x^3}} = 3\sqrt\frac{x^7}{x^3} = 3\sqrt\frac{x^3x^4}{x^3} = 3\sqrt{x^4} = 3x^2 \]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome